2. Calculate the maximum power transferred to the resistance R 5 Q RL 0.2v 10 82 4 V
Answer #1

7 Given circuit. RL 52 + Elon อ 2.0 0.20 - For finding maximum power transfered to network resistance RL, we have to find theAt node a. 20 :0.Qui TV-2-460 Vo = 0.212:46) Yo = 0.492 V2 - 4.92 V => Uth = vi-V2 = 2.46-492 = -2.460 To find thi we haveThevenins equivalent circuit , Въ 132 ŹRL=1332 2.46V maximum For Power P de circuit resistive load, according to tears for th